[Koji Suresh Kiwamu]
Suresh modified the MC board to have +5V offset at the output. (To be reported in the separated elog)
The MC lock has not been obtained at this point. An investigation revealed that there was very small (~5mVpp) PDH signal.
Kiwamu removed his triple resonant adapter and put the 50Ohm termination insted.
This restored the signal level normal although this changed the demodulation phase about 20deg.
We left the demodulation phase as it is because this is a temporary setup and the loss of the signal is not significant.
Now the MC is steadily locked with the single super boost.